A yacht used a cylindrical winder to unwind a cable to release an anchor. Using energy methods,



what is the cable’s acceleration? Explain your answer

1
Expert's answer
2021-12-20T10:31:23-0500

Find the angular velocity:


"mgh=\frac{1}{2}I\omega^2,\\\space\\
\omega=\sqrt\frac{2mgh}{I}."


The final velocity of the anchor:


"v=\omega R=R\sqrt\frac{2mgh}{I}."

The cable's acceleration (for the cable of length L):


"a=\frac{v^2}{2L}=\frac{2mghR^2}{2LI}."